OpenStudy (anonymous):

Find all zeros of the function f(x)=8(x-4)^4(9x+2)(x^2-9)and their multilpicties.

OpenStudy (anonymous):

f(x)=8(x-4)^4(9x+2)(x^2-9) =8(x-4)^4(9x+2)(x-3)(x+3) So the zeros are x=4 ; multiplicity 4 x=-2/9 multiplicity 1 x=1 multiplicity 1 x=-1 multiplicity 1

OpenStudy (anonymous):

Do you mean 3 and -3 instead of 1?

OpenStudy (anonymous):

Yes, he does.

OpenStudy (anonymous):

ty!